OpenStudy (anonymous):

A water park charges $5 for entry into the park and an additional $2 for each of the big waterslides. Steven spent $17 on his visit to the water park. Which equation can be used to find the number of big waterslides, x, that Steven went on?

OpenStudy (anonymous):

5 + 2x = 17 5x + 2 = 17 5x + 2x = 17 5 + 2 = 17x

OpenStudy (anonymous):

oh okay that would be 1st one :)